# Graphon Quantum Filtering systems

### v1

@@ 7 Conclusion and perspectives
We have extended the framework for studying continuously observed systems composed of a large number of interacting quantum particles with heterogeneous interactions, leading to graphon-based system of quantum filtering equations. We established a convergence result in the case of block-wise interactions with bosonic by class system. A natural direction for future research is to further investigate the question of mean-field limit for a general non-exchangeable system. Additionally, while our analysis assumes a sequence of dense graphs, one can relax this hypothesis to include moderately sparse (not-so-dense) graphs, using similar arguments as in [9]. Moreover, all results in this paper can be directly adapted for counting measurement case when the coupling operator $L$ is unitary, i.e. $L^{\dagger}=L^{-1}$. For more general coupling operator, the mean-field limit for $N$-particle system is still an open question, see the discussion in [41, remarks 5 ].
@@ Appendix A Exchangeable quantum particles
where $\underline{P}_{N}$ is a probability measure on $\mathfrak{E}^{N}$, with $\mathfrak{E}^{N}$ being the $N-$fold Cartesian product of a measurable space $\mathfrak{E}$.
@@
Appendix B Link between linear and non-linear form
In this section, we provide a strong solution to the linear system (4.4) and use it to construct a weak solution of the non-linear system (4.3).
Proposition 3 (Linear system SDE on Liouville space).
Let $L\in\mathcal{B}(\mathbb{H})$ and $(I\times\Omega,\mathcal{V},\mathcal{Q})$ be a Fubini extended filtered probability space carrying a family of e.p.i. Wiener processes $\mathbf{Y}=\big{\{}Y^{u}:u\in I\big{\}}$. For every $u\in I,$ let $(H_{t,u})_{t\geq 0}$ be a measurable mapping $t\mapsto H_{t,u}\in\mathcal{B}(\mathbb{H})$ with
$\sup_{0\leq s\leq T,\;u\in I}\|H_{s,u}\|<\infty,$
and assume $H_{t,u}=H_{t,u}^{\dagger}$. Fix a measurement efficiency $\eta\in(0,1]$ and an initial density operator $\rho_{0}\in\mathcal{S}(\mathbb{H})$. Consider, for each $u\in I$, the unnormalized Liouville–space valued stochastic differential system :
@@ Appendix B Link between linear and non-linear form
$\varrho_{0,u}$ $=\rho_{0}\in\mathcal{S}(\mathbb{H}).$
Then the system has a unique global solution that preserves positivity and
@@
First, existence and uniqueness are addressed directly via standard arguments on stochastic differential equations in infinite-dimensional Hilbert spaces [31].
@@
The process $(\bar{\varrho}_{t,u})_{t\geq 0}$ restricted to $\mathcal{V}_{t}^{Y^{u}}$ is equivalent to $({\varrho}_{t,u})_{t\geq 0}$, i.e.,
$\mathbb{E}_{\mathcal{Q}}[\bar{\varrho}_{t,u}|\mathcal{V}_{t}^{Y^{u}}]\sim{\varrho}_{t,u},\;\forall t\geq 0.$
Since $\bar{\varrho}_{0}={\varrho}_{0}=\rho_{0}\in\mathcal{S}(\mathbb{H})$, there exist a sequence of element in $\mathbb{H}$, $\{(\varphi_{k})_{k\geq 1}\}$ such that [^2] Every quantum state can be written as convex combination of pure states.
$\rho_{0}$ $=\sum_{k=1}^{\infty}\lambda_{k}\ket{\varphi_{0,k}}\bra{\varphi_{0,k}},$
@@
Consider a finite-dimensional truncation of the Hilbert space $\mathbb{H}_{d}:=P_{d}\mathbb{H}$ of dimension $d$, and a finite dimensional truncation $\rho_{0}^{d}$ of $\rho_{0}$
$\rho_{0}^{d}=\sum_{k=1}^{d}\lambda_{k}^{d}\ket{\varphi_{0,k}^{d}}\bra{\varphi_{0,k}^{d}}\longrightarrow\rho_{0}=\sum_{k=1}^{\infty}\lambda_{k}\ket{\varphi_{0,k}}\bra{\varphi_{0,k}}.$ | .
@@
$\mathrm{d}{\bar{\varrho}_{t,u}^{d}}$ $=-\mathrm{i}[H_{t},{\bar{\varrho}_{t,u}^{d}}]\mathrm{d}t+\Big{(}L{\bar{\varrho}_{t,u}^{d}}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,{\bar{\varrho}_{t,u}^{d}}\}\Big{)}\mathrm{d}t+\Big{(}L{\bar{\varrho}_{t,u}^{d}}+{\bar{\varrho}_{t,u}^{d}}L^{\dagger}\Big{)}\mathrm{d}X_{t}^{u}$
${\bar{\varrho}_{0,u}^{d}}$ $=\rho_{0}^{d}=\sum_{k=1}^{d}\lambda_{k}^{d}\ket{\varphi_{0,k}^{d}}\bra{\varphi_{0,k}^{d}}\in\mathcal{S}(\mathbb{H}_{d})$ (B.1)
@@
[…]
By Itô formula, |$\ket{\varphi_{t,u,k}^{d}}\bra{\varphi_{t,u,k}^{d}}$ | is a solution of (B.1). By a classical result [6, Proposition 2.8], there exists an invertible propagator $(\mathcal{U}_{t,u,k}^{d})_{t\geq 0}$ such that
@@
The invertibility of $\mathcal{U}_{t,u,k}^{d}$ and $\rho_{0}^{d}\neq 0$ guarantees the positivitiy of $\bar{\varrho}^{d}_{t,u}$ for all $t\in[0,T]$. In order to show the positivity of $(\bar{\varrho}_{t})_{t\geq 0}$, it is sufficient to extract a sequence $(\bar{\varrho}_{t}^{d})_{t\geq 0}$ that converges in the weak-sense i.e., for some sequence $(d_{j})_{j\geq 0}$,
@@
$\sup_{d}\mathbb{E}\Big{[}\sup_{0\leq t\leq T}|\bar{\varrho}_{t,u}^{d}|\Big{]}<\infty.$ (B.3)
To show (B.3), let us denote $\mathrm{t}\mathrm{r}(\bar{\varrho}_{t,u}^{d})=\sum_{k=1}^{d}\lambda_{k}^{d}\|\varphi_{t,u,k}^{d}\|^{2}$ and $\quad M_{t,k}^{d,u}:=\|\varphi_{t,u,k}^{d}\|^{2}$. From Itô’s formula,
@@
$\mathbb{E}\Big{[}\sup_{0\leq s\leq T}\|\varphi_{s,u,k}^{d}\|^{2}\Big{]}\leq C<\infty,$
and hence,
$\sup_{d}\mathbb{E}_{\mathcal{Q}}\bigl{[}\sup_{0\leq t\leq T}\mathrm{t}\mathrm{r}(\bar{\varrho}_{t,u}^{d})\bigr{]}\leq C.$
@@
Then $\bar{\varrho}^{d_{j}}_{\cdot,u}$ converges weakly to a positive operator valued process $\bar{\varrho}_{\cdot,u}$. Therefore, $\bar{\varrho}_{t,u}$ is positive, and hence so is $\varrho_{t,u}$. This completes the proof. ∎
We now state a version of Girsanov’s theorem on Fubini extension spaces to establish a link between the linear and nonlinear equations.
Lemma 5 (Girsanov theorem on Fubini extension).
Under a Fubini extension probability space $(I\times\Omega,\mathcal{I}\boxtimes\mathcal{F},\lambda\boxtimes\mathbb{Q})=(I\times\Omega,\mathcal{V},\mathcal{Q})$, we consider a family of essentially pairwise independent Wiener processes $\mathbf{Y}=\big{\{}Y^{u}:u\in I\big{\}}$, and $\mathcal{I}\boxtimes\mathcal{F}$-adapted process $(\theta_{t})_{t\geq 0}$, such that
$\int_{0}^{T}\|\theta_{t}\|^{2}_{L^{2}}\mathrm{d}t<\infty,\quad\text{i.e.,}\quad\int_{0}^{T}\Big{(}\int_{I}\int_{\Omega}|\theta_{t,u}(\omega)|^{2}\mathbb{Q}(\mathrm{d}\omega)\lambda(\mathrm{d}u)\Big{)}\mathrm{d}t<\infty.$
Define $\mathcal{E}_{t}=(\mathcal{E}_{t,u})_{u\in I}$ by
$\mathcal{E}_{t,u}:=\exp\Big{\{}\int_{0}^{t}\theta_{s,u}\mathrm{d}Y_{s}^{u}-\frac{1}{2}\int_{0}^{t}\theta_{s,u}^{2}\mathrm{d}s\Big{\}},$
such that $(\mathcal{E}_{t,u})_{t\geq 0}$ is a martingale for every $u\in I$. Then, under $\mathcal{P}$ with density $\mathcal{E}_{T}$ with respect to $\mathcal{Q}$, the collection $\mathbf{W}=\big{\{}W^{u}:u\in I\big{\}}$ forms a family of pairwise independent Wiener processes, where
$W_{t}^{u}=Y_{t}^{u}-\int_{0}^{t}\theta_{s,u}\mathrm{d}s.$
Proposition 4.
Let $(I\times\Omega,\mathcal{V},\mathcal{Q})$ be a fixed Fubini extended probability space and $\mathbf{Y}=\big{\{}Y^{u}:u\in I\big{\}}$ be a family of essentially pairwise independent Wiener processes. Suppose that $(\varrho_{t,u})_{t\geq 0}$ is a stochastic process valued on $\mathcal{L}_{2}(\mathbb{H})$, issued from the linear quantum stochastic master equation
@@
$\mathrm{d}\varrho_{t,u}$ $=-\mathrm{i}[H_{t,u},\varrho_{t,u}]\mathrm{d}t+\Big{(}L\varrho_{t,u}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\varrho_{t,u}\}\Big{)}\mathrm{d}t+\sqrt{\eta}\big{(}L\varrho_{t,u}+\varrho_{t,u}L^{\dagger}\big{)}\mathrm{d}Y_{t}^{u},$
$\varrho_{0,u}$ $=\rho_{0}\in\mathcal{S}(\mathbb{H}),$
@@
then the normalized process $\rho_{t,u}:=\frac{\varrho_{t,u}}{\mathrm{t}\mathrm{r}(\varrho_{t,u})}$ is a solution to the non-linear quantum stochastic master equation
@@
$+\sqrt{\eta}\Big{(}L\rho_{t,u}+\rho_{t,u}L^{\dagger}-\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\rho_{t,u}\Big{)}\mathrm{d}W_{t}^{u},$ (B.4)
where the family $\mathbf{W}=\big{\{}W^{u}:u\in I\big{\}}$ is a family of essentially pairwise independents Wiener processes on Fubini extension space $(I\times\Omega,\mathcal{V},\mathcal{P})$, with
$W_{t}^{u}:=Y_{t}^{u}-\sqrt{\eta}\int_{0}^{t}\mathrm{t}\mathrm{r}\Big{(}(L+L^{\dagger})\rho_{s,u}\Big{)}\mathrm{d}s,$
and the probability measure $\mathcal{P}$ has Radon-Nikodym derivative $\mathcal{E}_{t}:(u,\omega)\to\mathcal{E}_{t,u}(\omega)$ with respect to $\mathcal{Q}$, given by
$\frac{\mathrm{d}\mathcal{P}}{\mathrm{d}\mathcal{Q}}\Big{|}_{\mathcal{V}_{T}}$ $=\mathbf{\mathcal{E}}_{T},\quad{\mathcal{E}}_{t,{u}}:=\exp\Big{\{}\int_{0}^{t}\sqrt{\eta}\mathrm{t}\mathrm{r}\big{(}(L+L^{\dagger})\rho_{s,u}\big{)}\mathrm{d}Y_{s}^{u}-\frac{\eta}{2}\int_{0}^{t}\mathrm{t}\mathrm{r}^{2}\big{(}(L+L^{\dagger})\rho_{s,u}\big{)}\mathrm{d}s\Big{\}}.$
@@
We proceed ”à la Itô” to derive the nonlinear equation (B.4) from the linear dynamics:
@@
$\quad-\sqrt{\eta}\Big{(}L\rho_{t,u}+\rho_{t,u}L^{\dagger}-\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\rho_{t}\Big{)}\sqrt{\eta}\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\mathrm{d}t$
$=$ $\mathcal{L}_{t,u}(\rho_{t,u})\mathrm{d}t+\sqrt{\eta}\Big{(}L\rho_{t,u}+\rho_{t}L^{\dagger}-\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\rho_{t,u}\Big{)}\underbrace{(\mathrm{d}Y_{t}^{u}-\sqrt{\eta}\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t})\mathrm{d}t)}_{\mathrm{d}W_{t}^{u}},$

### v3

@@ 7 Conclusion and perspectives
We have introduced a framework for studying continuously observed systems composed of a large number of interacting quantum particles with heterogeneous interactions, and showed that the resulting graphon-based system of quantum filtering equations is well-posed. We established a convergence result in the case of block-wise interactions with bosonic-by-class structure. A natural direction for future research is to further investigate the mean-field limit for general non-exchangeable systems.
In addition, while our analysis assumes a sequence of dense graphs, this hypothesis can be relaxed to include moderately sparse graphs, using similar arguments as in [11, 31]. This can be seen as a bond percolation problem, where the $N$-body Hamiltonian is modified as
$\mathbf{H}^{N}$ $:=\sum_{q\in V_{N}}\mathbf{\tilde{H}}_{q}+\frac{1}{N\mathfrak{p}_{N}}\sum_{q\in V_{N}}\sum_{p>q}\xi^{N}_{pq}\mathbf{A}_{pq},$
with percolation parameter $\mathfrak{p}_{N}$ a sequence of positive real numbers that controls the sparsity of the graph, and such that $\mathfrak{p}_{N}\to 0$ while $N\mathfrak{p}_{N}\to\infty$ as $N\to\infty$.
Moreover, all results in this paper can be directly adapted to the counting measurement case (see the corresponding equation in the introduction) when the coupling operator $L$ is unitary, i.e., $L^{\dagger}=L^{-1}.$ For more general coupling operator, however, the mean-field limit for $N$-particle system remains an open problem, see the discussion in [45, Remark 5]. We leave this and related questions for future work.
@@ Appendix A Exchangeable quantum particles
where $\underline{P}_{N}$ is a probability measure on $\mathfrak{E}^{N}$, with $\mathfrak{E}^{N}$ being the NN-fold Cartesian product of a measurable space $\mathfrak{E}$.
@@
Appendix B Link between the linear and non-linear forms
In this section, we provide a strong solution to the linear system (10) and use it to construct a weak solution of the non-linear system (9).
We start by stating an estimate lemma, first derived in [48], which is useful for ensuring well-posedness in Liouville space. For completeness, we provide here a shorter proof that avoids the approximation arguments used in the original reference.
Lemma B.1 (Estimate lemma).
Let $A$ be a bounded self-adjoint operator and $B$ a bounded operator. Then:
  • (i)
    $2|\mathrm{t}\mathrm{r}(ABAB^{\dagger})|\leq\mathrm{t}\mathrm{r}(A^{2}(BB^{\dagger}+B^{\dagger}B))$,
  • (ii)
    $|\mathrm{t}\mathrm{r}(AB^{\dagger}AB^{\dagger}+ABAB)|\leq\mathrm{t}\mathrm{r}(A^{2}(BB^{\dagger}+B^{\dagger}B))$.
Proof.
For the first estimate, applying the Cauchy-Schwarz inequality, we get
$\big|\mathrm{t}\mathrm{r}(B^{\dagger}ABA)\big|\leq\sqrt{\mathrm{t}\mathrm{r}(B^{\dagger}A^{2}B)}\sqrt{\mathrm{t}\mathrm{r}(BA^{2}B^{\dagger})}=\sqrt{\mathrm{t}\mathrm{r}(A^{2}BB^{\dagger})}\sqrt{\mathrm{t}\mathrm{r}(A^{2}B^{\dagger}B)}.$
[…]
$\big|\mathrm{t}\mathrm{r}(ABAB^{\dagger})\big|\leq\tfrac{1}{2}\Bigg(\mathrm{t}\mathrm{r}(A^{2}BB^{\dagger})+\mathrm{t}\mathrm{r}(A^{2}B^{\dagger}B)\bigg).$
For the second estimate, use the triangular inequality:
$\Big|\mathrm{t}\mathrm{r}(AB^{\dagger}AB^{\dagger}+ABAB)\Big|$ $\leq\Big|\mathrm{t}\mathrm{r}(AB^{\dagger}AB^{\dagger})\Big|+\Big|\mathrm{t}\mathrm{r}(ABAB)\Big|.$
By the Cauchy-Schwarz inequality,
$\Big|\mathrm{t}\mathrm{r}(ABAB)\Big|=\Big|\mathrm{t}\mathrm{r}(AB(B^{\dagger}A)^{\dagger})\Big|\leq\sqrt{\mathrm{t}\mathrm{r}(ABB^{\dagger}A)}\sqrt{\mathrm{t}\mathrm{r}(B^{\dagger}A^{2}B)}=\mathrm{t}\mathrm{r}(A^{2}BB^{\dagger}),$
and similarly,
$\Big|\mathrm{t}\mathrm{r}(AB^{\dagger}AB^{\dagger})\Big|\leq\mathrm{t}\mathrm{r}(A^{2}B^{\dagger}B).$
Hence,
$\Big|\mathrm{t}\mathrm{r}(AB^{\dagger}AB^{\dagger}+ABAB)\Big|\leq\mathrm{t}\mathrm{r}(A^{2}(B^{\dagger}B+BB^{\dagger})).$
Throughout this part we work on a Fubini extended probability space $(I\times\Omega,\mathcal{V},\mathcal{Q})$ of the usual product space $(I\times\Omega,\mathcal{I}\otimes\mathcal{F},\lambda\otimes\mathbb{Q})$. This space carries a family of e.p.i. Wiener processes $\mathbf{Y}=\{Y^{u}:u\in I\},$ and we fix a bounded operator $L\in\mathcal{B}(\mathbb{H})$.
Under this standing framework, we state the following results.
Proposition B.2 (Linear system SDE on Liouville space).
Let $\mathbf{H}=\{H_{u}:u\in I\}$ be a measurable family such that
$t\mapsto H_{t,u}\in\mathcal{B}(\mathbb{H}),\qquad\sup_{0\leq s\leq T,\,u\in I}\|H_{s,u}\|<\infty,$
and assume $H_{t,u}=H_{t,u}^{\dagger}$ for all $(t,u)\in[0,T]\times I$.
Fix $\eta\in(0,1]$ and an initial density operator $\rho_{0}\in\mathcal{S}(\mathbb{H})$. Consider the Liouville–space valued stochastic differential equation
$\mathrm{d}\mathrm{R}_{t}$ $=\mathcal{L}_{t}(\mathrm{R}_{t})\mathrm{d}t+\sqrt{\eta}\mathcal{R}(\mathrm{R}_{t})\mathrm{d}\mathbf{Y}_{t},$ (17)
$\mathrm{R}_{0}$ $=\rho_{0}\in\mathcal{B}_{2}(\mathbb{H}).$
The drift $\mathcal{L}_{t}$ is defined by
$\mathcal{L}_{t}(R)=-\mathrm{i}[\mathbf{H}_{t},R]+\big(LRL^{\dagger}-\frac{1}{2}\{L^{\dagger}L,R\}\big),$
and the diffusion $\mathcal{R}$ by
$\mathcal{R}(R)=LR+R{L^{\dagger}}.$
Coordinatewise, the equation reads
@@ Appendix B Link between the linear and non-linear forms
$\varrho_{0,u}$ $=\rho_{0}\in\mathcal{B}_{2}(\mathbb{H}).$
Then the equation has a unique global solution. If $\rho_{0}\in\mathcal{S}(\mathbb{H})$, then positivity is preserved and
@@
First, the existence and uniqueness are addressed directly via standard arguments on stochastic differential equations in infinite-dimensional Hilbert spaces [32].
@@
The process $(\bar{\varrho}_{t,u})_{t\geq 0}$ restricted to $\sigma({Y^{u}_{t}})$ is equivalent to $({\varrho}_{t,u})_{t\geq 0}$, i.e.,
$\mathbb{E}_{\mathbb{Q}}[\bar{\varrho}_{t,u}|\sigma({Y}^{u}_{t})]\sim{\varrho}_{t,u},\;\forall t\geq 0.$
Since $\bar{\varrho}_{0}={\varrho}_{0}=\rho_{0}\in\mathcal{S}(\mathbb{H})$, there exists a sequence of elements in $\mathbb{H}$, $\{(\varphi_{k})_{k\geq 1}\}$ such that  [^2] Every quantum state can be written as a convex combination of pure states.
$\rho_{0}$ $|,\displaystyle=\sum_{k=1}^{\infty}\lambda_{k}\ket{\varphi_{0,k}}\bra{\varphi_{0,k}}$,
@@
Consider a finite-dimensional truncation of the Hilbert space $\mathbb{H}_{d}:=P_{d}\mathbb{H}$ of dimension $d$, and a finite dimensional truncation $\rho_{0}^{d}$ of $\rho_{0}$, given by
[…]
@@
$\mathrm{d}{\bar{\varrho}_{t,u}^{d}}$ $=-\mathrm{i}[H_{t},{\bar{\varrho}_{t,u}^{d}}]\mathrm{d}t+\Big(L{\bar{\varrho}_{t,u}^{d}}L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,{\bar{\varrho}_{t,u}^{d}}\}\Big)\mathrm{d}t+\Big(L{\bar{\varrho}_{t,u}^{d}}+{\bar{\varrho}_{t,u}^{d}}L^{\dagger}\Big)\mathrm{d}X_{t}^{u},$
[…]
@@
$\mathrm{d}\varphi_{t,u,k}^{d}$ $=-\mathrm{i}H_{t,u}\varphi_{t,u,k}^{d}\mathrm{d}t-\frac{1}{2}L^{\dagger}L\varphi_{t,u,k}^{d}\mathrm{d}t+L\varphi_{t,u,k}^{d}\mathrm{d}X_{t}^{u},\;1\leq k\leq d.$ (19)
By Itô formula, $|\ket{\varphi_{t,u,k}^{d}}\bra{\varphi_{t,u,k}^{d}}$ is a solution of (18). By a classical result [8, Proposition 2.8], there exists an invertible propagator $(\mathcal{U}_{t,u,k}^{d})_{t\geq 0}$ such that
@@
The invertibility of $\mathcal{U}_{t,u,k}^{d}$ and $\rho_{0}^{d}\neq 0$ guarantee the positivity of $\bar{\varrho}^{d}_{t,u}$ for all $t\in[0,T]$. In order to show the positivity of $(\bar{\varrho}_{t})_{t\geq 0}$, it is sufficient to extract a sequence $(\bar{\varrho}_{t}^{d})_{t\geq 0}$ that converges in the weak-sense, i.e., for some sequence $(d_{j})_{j\geq 0}$,
@@
$\sup_{d}\mathbb{E}\Big[\sup_{0\leq t\leq T}|\bar{\varrho}_{t,u}^{d}|\Big]<\infty.$ (20)
To show (20), let us denote $\mathrm{t}\mathrm{r}(\bar{\varrho}_{t,u}^{d})=\sum_{k=1}^{d}\lambda_{k}^{d}\|\varphi_{t,u,k}^{d}\|^{2}$ and $\quad M_{t,k}^{d,u}:=\|\varphi_{t,u,k}^{d}\|^{2}$. From Itô’s formula,
@@
$\mathbb{E}_{\mathbb{Q}}\Big[\sup_{0\leq s\leq T}\|\varphi_{s,u,k}^{d}\|^{2}\Big]\leq C<\infty,$
and therefore
$\sup_{d}\mathbb{E}_{\mathbb{Q}}\bigl[\sup_{0\leq t\leq T}\mathrm{t}\mathrm{r}(\bar{\varrho}_{t,u}^{d})\bigr]\leq C.$
@@
Then $\bar{\varrho}^{d_{j}}_{\cdot,u}$ converges weakly to a positive operator valued process $\bar{\varrho}_{\cdot,u}$. Hence, $\bar{\varrho}_{t,u}$ is positive, and consequently so is $\varrho_{t,u}$.
In order to conclude, we have to show that $R_{t}\in L^{2}_{\mathcal{L}_{2}(\mathbb{H})}(I\times\Omega,\mathcal{Q})$. We proceed “à la Grönwall”:
$\mathrm{d}\mathrm{t}\mathrm{r}(\varrho_{t,u}^{2})$ $=\mathrm{t}\mathrm{r}\big(-\mathrm{i}[H_{t,u},\varrho_{t,u}]\varrho_{t,u}-\mathrm{i}\varrho_{t,u}[H_{t,u},\varrho]\big)\mathrm{d}t$
$+\mathrm{t}\mathrm{r}\big((L\varrho_{t,u}L^{\dagger}$ $-\frac{1}{2}L^{\dagger}L\varrho_{t,u}-\frac{1}{2}\varrho_{t,u}L^{\dagger}L)\varrho_{t,u}+\varrho_{t,u}(L\varrho_{t,u}L^{\dagger}-\frac{1}{2}L^{\dagger}L\varrho_{t,u}-\frac{1}{2}\varrho_{t,u}L^{\dagger}L)\big)\mathrm{d}t,$
$+\eta\mathrm{t}\mathrm{r}\big($ $(L\varrho_{t,u}+\varrho_{t,u}L^{\dagger})(L\varrho_{t,u}+\varrho_{t,u}L^{\dagger})\big)\mathrm{d}t+(\dots)\mathrm{d}Y_{t}^{u}$
@@
$\frac{\mathrm{d}\mathbb{E}_{\mathbb{Q}}[\mathrm{t}\mathrm{r}(\varrho_{t,u}^{2})]}{\mathrm{d}t}$ $=\!\mathbb{E}_{\mathbb{Q}}\Biggl[2\mathrm{t}\mathrm{r}\bigl(\varrho_{t,u}L\varrho_{t,u}L^{\dagger}\!-\!\varrho_{t,u}^{2}L^{\dagger}L\bigr)+\eta\mathrm{t}\mathrm{r}\bigl(L\varrho_{t,u}L\varrho_{t,u}+L^{\dagger}\varrho_{t,u}L^{\dagger}\varrho_{t,u}+2L^{\dagger}L\varrho_{t,u}^{2}\bigr)\Biggr],$
$\frac{\mathrm{d}\mathbb{E}_{\mathbb{Q}}[\mathrm{t}\mathrm{r}(\varrho_{t,u}^{2})]}{\mathrm{d}t}$ $=\!\mathbb{E}_{\mathbb{Q}}\!\Biggl[\mathrm{t}\mathrm{r}\Bigl(2\varrho_{t,u}L\varrho_{t,u}L^{\dagger}+\eta\bigl(L\varrho_{t,u}L\varrho_{t,u}+L^{\dagger}\varrho_{t,u}L^{\dagger}\varrho_{t,u}\bigr)\Bigr)-2(1-\eta)\mathrm{t}\mathrm{r}\bigl(L^{\dagger}L\varrho_{t,u}^{2}\bigr)\Biggr],$
and using Lemma B.1,
$\mathrm{d}\mathbb{E}_{\mathbb{Q}}[\mathrm{t}\mathrm{r}(\varrho_{t,u}^{2})]$ $\leq\mathbb{E}_{\mathbb{Q}}\Biggl[(1+\eta)\mathrm{t}\mathrm{r}\bigl(\varrho_{t,u}^{2}\bigl(L^{\dagger}{L}+{L}{L}^{\dagger}\bigr)\bigr)+2(1-\eta)\mathrm{t}\mathrm{r}\bigl(L^{\dagger}L\varrho_{t,u}^{2}\bigr)\Biggr]\mathrm{d}t$
$\leq 2(1+\eta)\|L\|^{2}\mathbb{E}_{\mathbb{Q}}\bigl[\mathrm{t}\mathrm{r}(\varrho_{t,u}^{2})\bigr]\mathrm{d}t+4(1-\eta)\|L\|^{2}\mathbb{E}_{\mathbb{Q}}\bigl[\mathrm{t}\mathrm{r}(\varrho_{t,u}^{2})\bigr]\mathrm{d}t$
$\leq(6-2\eta)\|L\|^{2}\mathbb{E}_{\mathbb{Q}}\bigl[\mathrm{t}\mathrm{r}(\varrho_{t,u}^{2})\bigr]\mathrm{d}t.$
Thanks to Grönwall’s inequality,
$\mathbb{E}_{\mathbb{Q}}[\mathrm{t}\mathrm{r}(\rho_{t,u}^{2})]$ $\leq e^{\big((6-2\eta)\|L\|^{2}\big)t}.$
Then,
$\mathbb{E}_{\mathcal{Q}}[\|\mathrm{R}_{t}\|^{2}_{2}]$ $=\int_{I\times\Omega}\|\mathrm{R}_{t}(u,\omega)\|^{2}_{2}\mathcal{Q}(\mathrm{d}u,\mathrm{d}\omega),$
$=\int_{I}\mathbb{E}_{\mathbb{Q}}[\mathrm{t}\mathrm{r}(\varrho_{t,u}^{2})]\lambda(\mathrm{d}u),$
$\leq e^{\big((6-2\eta)\|L\|^{2}\big)T}<\infty.$
This completes the proof. ∎
By the Girsanov theorem stated in Proposition 2.8, we establish a link between the linear and nonlinear equations.
Proposition B.3.
@@
Consider $\mathrm{R}_{t}=\{\varrho_{t,u}:u\in I\}$ a stochastic process valued on $\mathcal{L}_{2}(\mathbb{H})$, issued from the linear quantum stochastic master equation (17). Then the process $\bar{\mathrm{R}}_{t}=\{\rho_{t,u}:u\in I\}$, with $\rho_{t,u}:=\frac{\varrho_{t,u}}{\mathrm{t}\mathrm{r}(\varrho_{t,u})}$, is a solution to the non-linear quantum stochastic master equation
@@
$+\sqrt{\eta}\Big(L\rho_{t,u}+\rho_{t,u}L^{\dagger}-\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\rho_{t,u}\Big)\mathrm{d}W_{t}^{u},$ (21)
where the process $\mathbf{W}=\big\{W^{u}:u\in I\big\}$ is a Wiener process under the new probability measure $\tilde{\mathcal{P}}$, defined by the Girsanov transform
$\mathbf{W}_{t}:=\mathbf{Y}_{t}-\sqrt{\eta}\int_{0}^{t}\mathrm{t}\mathrm{r}\big((L+L^{\dagger})\bar{\mathrm{R}}_{s}\big)\mathrm{d}s.$
The measure $\tilde{\mathcal{P}}$ has Radon-Nikodym derivative $\mathcal{E}_{t}:(u,\omega)\to\mathcal{E}_{t,u}(\omega)$ with respect to $\mathcal{Q}$, given by
${\mathcal{E}}_{t,{u}}:=\exp\Big\{\int_{0}^{t}\sqrt{\eta}\mathrm{t}\mathrm{r}\big((L+L^{\dagger})\rho_{s,u}\big)\mathrm{d}Y_{s}^{u}-\frac{\eta}{2}\int_{0}^{t}\mathrm{t}\mathrm{r}^{2}\big((L+L^{\dagger})\rho_{s,u}\big)\mathrm{d}s\Big\}.$
@@
We proceed “à la Itô” to derive the nonlinear Equation (21) from the linear dynamics:
@@
$\quad-\sqrt{\eta}\Big(L\rho_{t,u}+\rho_{t,u}L^{\dagger}-\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\rho_{t,u}\Big)\sqrt{\eta}\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\mathrm{d}t$
$=$ $\mathcal{L}_{t,u}(\rho_{t,u})\mathrm{d}t+\sqrt{\eta}\Big(L\rho_{t,u}+\rho_{t}L^{\dagger}-\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\rho_{t,u}\Big)\underbrace{(\mathrm{d}Y_{t}^{u}-\sqrt{\eta}\mathrm{t}\mathrm{r}((L+L^{\dagger})\rho_{t,u})\mathrm{d}t)}_{\mathrm{d}W_{t}^{u}},$